### v1

# Moving towards informative and actionable social media research

Moving towards informative and actionable social media research
Joseph B. Bak-Coleman 1,2,3,4,5∗ Stephan Lewandowsky 6,7 Philipp Lorenz-Spreen 8,9∗ Arvind Narayanan 10,11 Amy Orben 12 Lisa Oswald8
1 Centre for the Advanced Study of Collective Behavior, University of Konstanz, Konstanz, Germany
2 Department of Collective Behavior, Max Planck Institute of Animal Behavior, Konstanz, Germany
3 Department of Biology, University of Konstanz, Konstanz, Germany
4 Santa Fe Institute, Santa Fe, NM, USA
5 Berkman Klein Center, Harvard University, Cambridge, MA, USA
6 School of Psychological Science, University of Bristol, Bristol, UK
7 Department of Psychology, University of Potsdam, Potsdam, Germany
8 Center Synergy of Systems and Center for Scalable Data Analytics and Artificial Intelligence, Dresden University of Technology, Dresden, Germany
9 Center for Adaptive Rationality, Max Planck Institute for Human Development, Berlin, Germany
10 Center for Information Technology Policy, Princeton University, Princeton, NJ, USA
11 Department of Computer Science, Princeton University, Princeton, NJ, USA
12 MRC Cognition and Brain Sciences Unit, University of Cambridge, Cambridge, UK
To whom correspondence should be addressed; E-mail: jbakcoleman@gmail.com, philipp.lorenz-spreen@tu-dresden.de
Social media is nearly ubiquitous in modern life, and concerns have been raised about its putative societal impacts, ranging from undermining mental health and exacerbating polarization to fomenting violence and disrupting democracy. Despite extensive research, consensus on these effects remains elusive, with observational studies often highlighting concerns while randomized controlled trials (RCTs) yield conflicting or null findings. This review examines how the complexity inherent in social systems can account for such discrepancies, emphasizing that emergent societal and long-term outcomes cannot be readily inferred from individual-level effects. In complex systems, such as social networks, feedback loops, hysteresis, multi-scale dynamics, and non-linearity limit the utility of approaches for assessing causality that are otherwise robust in simpler contexts. Revisiting large-scale experiments, we explore how null or conflicting findings may reflect these complexities rather than a true absence of effects. Even in cases where the methods are appropriate, assessing the net impacts of social media provides little actionable insight given that eliminating social media is not a realistic option for whole populations. We argue that progress will require a complexity-minded approach focused on specific design choices of online platforms that triangulates experimental, observational and theoretical methods.
Introduction
Social media have permeated most societies, impacting citizens lives around the world. Nearly 5 billion people across the globe are using social media, and penetration rates exceeding 80% in northern Europe, 70% in North-, 67% in South-America and reaching 75% in Eastern Asia, for example  (?). The defining attribute of social media platforms is that users not only consume but also create and distribute content. In addition, social media are characterized by some degree of algorithmic curation and in most cases also an advertisement-driven business model  (?).
Accompanying the pervasive use of social media is widespread concern about their adverse effects on society, ranging from undermining public health measures  (?) and mental health  (?, ?) to exacerbating polarization  (?), fomenting extremism  (?), and disrupting democracy  (?). A plethora of research has been seeking to address those concerns; Google Scholar returns more than 800,000 publications since 2020 in response to the query “social media” (Search on 16 February 2024). At least tacitly, this substantial research endeavor is concerned with generating a consensus causal understanding of the relationships between social media use and the outcomes of interest. Absent such an understanding, any necessary corrective action (including national or international regulation and intervention) may remain ineffective.
Unfortunately, most efforts trying to understand the impact of social media on large-scale societal changes —to ultimately judge whether they warrant corrective action— persist in the liminal space between compelling patterns or individual case studies of harm on one hand, and consensus causal estimates on the other. For example, a recent systematic review of the literature of social media’s impact on political behavior identified almost 500 papers  (?) and revealed a multi-dimensionality of associations. Many reviewed papers reported positive associations between social-media use and political participation and knowledge, and at least as many papers reported associations between social media use and growing polarization and success of populist parties, which point in a detrimental direction for liberal democracies.
Since then, most researcher agree to the correlational observations that were reported by Lorenz-Spreen et al., no strong consensus on the causal impacts and mechanisms of social media on society and democracy has emerged. Scientists remain divided about whether polarization is caused, amplified or just made visible by social media platforms  (?) and how or even whether that causes harm  (?). Similar discussions are also had regarding the impact of social media on mental health, violence and other societal outcomes of interest. Specifically, due to the lack of agreement on causal mechanisms, there are few (if any) concrete suggestions for policies or regulations that should be implemented to reduce harms and promote positive individual and societal outcomes in our digital age.
The inability to reach scientific consensus and provide actionable insight hints at the existence of deeper problems in the research endeavor that require resolution before progress can occur. Here, we argue that there are two key barriers to an effective science of social media. First, social media constitute complex systems that comprise ensembles of individual elements that interact with each other in specific ways, thereby forming higher-level order and emergent phenomena on the collective level (?, ?). However, scientists tend to use techniques such as randomized controlled trials (RCTs) that are well suited to simpler, socially-independent individual outcomes but ill-suited to establishing causality for complex systems outcomes (e.g., societal polarization), particularly in the context of social media. This leads to long-winded debate, as bias, confounds, and model assumption failures may drive observed effects more than the underlying phenomena of interest—a general challenge recently reviewed in the social sciences more broadly by Bailey et al  (?). Second, even when empirical circumstances allow answers to certain causal questions, various studies tend to produce only broad insights (e.g., about average effects of deactivation) that are not directly actionable by technologists and regulators.
To overcome those barriers, we propose (1) a complex systems approach to study social media’s impacts on society, which is based on the triangulation of mechanistic models, experiments suitable for estimating causal effects at a relevant scale, as well as descriptive work and observational causal inferences at levels of systemic complexity. (2) Furthermore, we suggest to focus research questions on more detailed understanding of relationships between platform affordances and behaviors or attitudes.
Research on Social Media’s Impacts
Research on social media’s impacts routinely uncovers associations suggesting links between platform use and broader societal or individual outcomes. For example, consistent correlational evidence has been identified across several dimensions, including between social media use and political participation, polarization, trust in political institutions, body image and well-being  (?, ?, ?). Aiming to gain causal insight into these associations, some studies have used quasi-experimental instrumental-variable approaches  (?) in which predictable variation of local internet adoption or social media use caused by a third variable (the “instrument”) has been used to estimate the causal impact on trust  (?, ?), political participation  (?, ?), hate crimes  (?, ?), or populist support and polarization  (?, ?), as well as mental health  (?, ?). The results from those naturalistic quasi-experiments yield a relatively consistent and concerning pattern, with social media being shown to increase hate crimes, lower trust, or increase polarization and support for populists (for a more detailed overview, see  (?)).
An alternative approach involves large-scale field experiments in which individuals are placed into randomized controlled trials (RCTs). Treatments are often constrained by limitations in independent researchers’ abilities to alter platform experiences and functionality. Examples range from participants following or subscribing to political accounts on Twitter/X and Facebook  (?, ?) through deactivation of Facebook for a month or a week  (?, ?, ?) to setting specific news websites as the starting page of a browser  (?), browsing allocated websites of a specific valence  (?) or implementing a suite of different approaches to reduce digital distractions  (?). More recent studies, conducted in collaboration with Meta, were able to go beyond these typical limitations, actively altering news-feed algorithms  (?, ?, ?). Another, recent study used a browser add-on to reduce toxic content within the feed and without direct collaboration with the platforms  (?).
Notably, results from those RCTs rarely seem to produce consensus in the social media context, (?) with the resulting findings conflicting both across RCTs as well as with inferences drawn from observational or quasi-experimental approaches. For example, two studies in which the treatment involved subscriptions/exposure to specific content affected polarization in opposite directions  (?, ?). A similar divergent pattern was observed between the two deactivation experiments of Facebook  (?, ?), where one study found reduced polarization from the deactivation the other found increased animosity against outgroups from the same treatment. In some cases, RCTs yield effects that are consistent with concerns raised from observational causality—yet effect sizes are small and may fail to achieve statistical significance. For example, in the most recent and largest deactivation experiment of Facebook and Instagram, affective polarization was slightly reduced and voting choice appeared altered  (?). However, this finding was judged not to be significant after application of a correction for multiple comparisons.
To further complicate the matter, outcomes from RCTs can differ substantially depending on what is measured. Results from interventions are often the most conflicting or weak when examining distal outcomes, such as polarization. In contrast, the impact of specific interventions on more proximal outcomes, such as exposure to harmful content, appear to be more stable. For example, the Meta collaboration generated consistent evidence that the type of content that individual people are exposed to is impacted by platform design and affordances, such as platform-mediated amplification of co-partisan or uncivil content  (?, ?, ?). On the other hand, the same studies showed that those changes had virtually no downstream consequences on outcomes of particular interest such as political attitudes  (?, ?, ?). As a result, there seems to be a missing link, whereby RCTs find consistent evidence with observational approaches for proximal outcome measures on the level of individual content exposure—yet effects on people’s distal behaviors and attitudes are conflicting or minimal. This creates a challenge for those who seek to attribute causal responsibility to platform algorithms for distal or societal-level outcomes such as polarization.
We argue that this apparent gap between individual-level outcomes (e.g., content exposure, engagement, sharing behaviour) and societal outcomes (e.g., polarization, out-group attitudes) arises because social media are inherently complex systems in which production and consumption of content are entangled in feedback loops that often operate across different time scales and levels of organizational complexity. Failing to account for such complexity when designing RCTs may compromise a causal interpretation of the results. Further, by casting aside these crucial attributes of complex systems, researchers have frequently focused on research questions that are in principle unanswerable or unactionable. If these questions are instead approached within a complex-systems perspective and associated methods, we will be one step closer to research that can lead to actionable policy, regulation, or design changes.
Causality in complex systems
There are certain questions about social media that are no longer answerable by direct observation. Just as we will never know what world we might inhabit now had the automobile not become the primary means of transport in America, we will never know how politics might have unfolded in the absence of Facebook or Twitter. The inability to examine counterfactual realities is known as the fundamental problem of causal inference, and is a widespread challenge across scientific research  (?).
Although causality cannot be directly measured in a strict sense, this has not stopped scientists from developing techniques to infer it indirectly  (?). Historically, the RCT is the principal workhorse in establishing causality. In an RCT, individual units of interest are randomly assigned to groups (e.g., treatment and control), and differences between those groups are evaluated. This design sidesteps the fundamental problem of causal inference by assuming that randomization, at least in expectation, eliminates systematic differences between the groups and yields an unbiased impact of treatment (but see  (?)). Early in their training, scientists are often taught that RCTs are the “gold-standard” of causality, which frequently results in misunderstanding of precisely how RCTs work, what they require, and what they yield  (?).
Most conspicuously, individual-level RCTs cannot reliably yield the estimand of interest for most complex social phenomenon, such as the causal effect of social media on collective processes such as polarization, extremism, mental health trends or vaccine uptake. To do so, one would have to randomize social media exposure at the scale of societies or (ideally) worlds. In lieu of this, RCTs are typically applied at the scale of individual users, under the buried assumption that societal effects over longer periods of time (which is often the actual estimand of interest) are equal to additive individual effects over much shorter periods of time (the estimated quantity). This assumption that causal impacts at societal scales should be observable at the individual level is tenuous, specifically for studying the societal impact of social media, which is a quintessential example of a complex system.
Complex systems are defined by the way in which interactions lead to emergent phenomena such that the behavior of the system cannot be understood as the sum of its parts (cf., ”More is Different”  (?)). Moreover, complex systems exhibit dynamics such that any estimated quantity is not independent of the point in time and historical state. Non-linearities introduced by such interactions produce feedback, phase-transitions, and hysteresis which preclude the notion of some time-invariant causal effect.
Figure 1: Illustrations of key properties of complex systems that make some RCTs particularly difficult to interpret. A Global outcomes do not linearly depend on the sum of their individual components, making conclusions from small groups to larger societies generally difficult. B More specifically, the outcomes depend on the history of the system, known as hysteresis, which makes conclusions about any reverse effects impossible. C Feedback loops that prevent conclusions about one variable without considering the value of another at a specific point in time. D Spillover effects in network-structures, known as violations of the “stable unit treatment value assumption” (SUTVA), prevent a clean separation of treatment and control conditions.
We highlight four key ways in which RCTs may yield misleading interpretations of causal effects both for effects that occur on longer time-scales or at larger scales of organizational complexity (see Figure 1.) We can illustrate those problems using an intuitive physical example. Consider concerns from observational evidence that power-grid blackouts coincide with increased demand for power in a world where the relationship between power demand and system failure is not known. Attempting to test this theory, an RCT is conducted on a power grid in which some substations are randomized to experience sudden spikes in demand. If the treatment group receiving increased load is small relative to the scale of the grid, the power grid will remain functional as the grid compensates for the localized demand. One would be forced to conclude that spikes of that magnitude have no effect on the power grid’s functionality. Conversely, if a critical number of substations had been exposed to the same spike in demand, blackouts would occur—including the control nodes for which no increase in demand occurred, because the blackouts would affect both types of nodes equally. Again, this would result in little difference between the groups, suggesting that blackouts occur independent of the condition (spike or not). Both of those interpretations would be incorrect because the estimated quantity from an individually targeted RCT is not the estimand of interest (also see Fig. 1A). In each of these cases, the outcome of interest cannot be isolated to individuals as the impacts of the treatment alter the system as a whole, in turn altering outcomes for both treatment and control groups alike.
The power grid example can also provide insight into the second problem, hysteresis (Figure 1B), which occurs when the relationship between inputs and the system state is not a simple function but depends on the past state of the system. Following failure to a power grid, an RCT relieving the load for some substations, but not others, would fail to bring any stations back online. Withdrawal studies rest on the assumption that the effect of the platform or feature can be inferred as the opposite of the measured withdrawal effect. We would be wrong to conclude from such a study that load did not cause the failure, merely because eliminating the load did not return function. This problem of interpretation besets all withdrawal studies, in which users in the treatment group are deactivating a platform, reducing use or even removing a feature  (?, ?, ?, ?).
In social systems, “path-dependence” is commonly used to describe hysteretic effects. To illustrate, suppose we want to evaluate the causal effect of Duolingo on language fluency, and we do so by randomly selecting long-time users to discontinue learning for a week. We are bound to find that users largely retain their fluency after a week of withdrawal of the platform. Of course, it would be incorrect to assume the lack of change in fluency reflects an inability of Duolingo to cause language learning. In general, there is no reason to believe that changes in beliefs, behaviors, attitudes (e.g., trust in institutions), or alterations of mental health acquired over years of time spent on a social media platform will rapidly or symmetrically revert upon brief (e.g., weeks) withdrawal, if such reversals are even possible or likely (also see Fig. 1B). Similar considerations apply to other social outcomes: For example, mental health might transition from one stable state (healthy) to another (unhealthy), but going back is not simply an additive feature of “undoing” the original trigger  (?). In general, we cannot reasonably assume that withdrawal from a social media platform produces equal and opposite effects as the initial exposure.
The third problem associated with interpreting individual-level RCTs is that the effects may unfold non-monotonically over time and depend on both the state of other variables in the system and the moment of observation (Figure 1C). For example, the feedback loop of addiction means that having a cigarette can lead to a relief (e.g., reduced tension or stress) in a long-term smoker. When such long-term smokers are asked to withdraw from smoking for two weeks, they are likely to become irritable (and gain weight, among other side effects of withdrawal). The clear health benefits of smoking cessation would only be observable over much larger time-scales. Under the reasoning common in social media studies RCTs, short-term studies of smoking cessation would incorrectly conclude that smoking is a safe and effective means of losing weight and reducing irritability (also see Fig. 1C). Likewise, relevant outcomes on social media may exhibit temporal dynamism—specifically recommendation algorithms that learn from user behaviour (?)—so we should anticipate that findings across experiments of differing lengths will conflict and short-term results may be of very limited value to extrapolate to the long term.
Finally, perhaps the core reason why RCTs are an inappropriate tool to study macro-outcomes on social media and power grids alike is that in both cases the individual units interact (Figure 1D). Formally, the problem here is that any RCT would violate the “stable unit treatment value assumption” (SUTVA), which requires that the treatment outcome of one experimental unit (e.g., a participant or power substation) is not affected by the treatment status of other units  (?). The violation of SUTVA impacts our powergrid RCT above, because whether or not a substation can continue running depends not only on its local load, but on the load which other directly and indirectly connected stations are experiencing and the power they are producing.
This problem also affects RCTs on social media. People are coupled through a myriad of on- and off-platform connections. An individual randomly exposed to an experimentally-altered newsfeed, for example by switching to a chronological sorting, will nonetheless experience algorithmic curation: When neighbors in their network share content they saw through algorithmic curation, the individual in the chronological condition is still indirectly subject to algorithmic sorting. Similarly, the mental well-being of a teenager taken off social media will not just depend on that experimental manipulation, but on whether their friends or parents also stop using these platforms.
Ultimately, violations of SUTVA can be avoided only if entire subnetworks are randomized into treatment and control. That is, all friends and family on social media must be assigned to the same treatment (or control) as the focal person of interest whose behaviour is recorded. To our knowledge, no existing RCTs fulfill that criterion.
In summary, any individually-focused RCT on social media that is not explicitly addressing the problems just presented—i.e., SUTVA violations, hysteresis, feedback loops and non-linearity inherent to complexity—is bound to be of limited value as clear causal identification of macro-level outcomes is not possible. The estimated magnitude and even direction of effects may be misleading, and there is no simple way to correct or adjust for such biases post hoc. It follows that any assertion that RCTs are interpretable or superior to observational approaches simply by virtue of being RCTs is not sustainable. The quality of any causal insight depends wholly on the assumptions underlying the causal inference and not the intrinsic properties of the approach. Very often the identification of a causal mechanism requires a multitude of methods, for example by triangulating it from observational, experimental and theoretical angles  (?). We explore examples of this problem using several recent studies in depth in the next section.
Example: RCTs on Facebook
A recent collaboration between Meta and independent researchers attempted to bridge the gap in our causal understanding of political outcomes algorithmic curation, using RCTs. In the experiments, some users were exposed to a number of interventions, such as receiving less politically-aligned content in their news feeds  (?), experiencing a reduced number of direct reshares from social contacts  (?), being switched to a chronologically sorted feed  (?), or deactivating the entire platform for 6 weeks  (?). Generally speaking, these interventions exhibited effects on exposure to content. For example, default platform features (e.g., algorithmic sorting) tended to expose users to more content that was ideologically aligned and uncivil but also less moderate sources, political content and untrustworthy sources compared to a reverse-chronological feed  (?).
At first glance, these findings are consistent with some concerns that engagement-optimized social media exacerbates polarization and foments hate via increased exposure to uncivil and like-minded content. However, the studies additionally conducted surveys to evaluate downstream effects on relevant attitudes, beliefs and behaviors. Here, effects were generally smaller, less consistent, and more ambiguous. For example, reducing exposure to content from like-minded sources did not decrease affective polarization or ideological extremity  (?). The authors asserted their estimated causal effects on individuals provided insight into the societal-scale estimand  (?).
We can reappraise those studies in light of the issues surrounding complexity raised above. First, as with our power grid and increased energy demand, there is little reason to believe that changes in collective state (e.g., polarization) from increased platform use or platform-feature exposure can be directly inferred from changes in exposure for an ensemble of individuals. Such changes involve not only the direct effects of exposure (or withdrawal), but also the changes following exposure to the dynamics across individuals (global outcomes, see Fig. 1A). Moreover, much like a power station remaining dark once demand is reduced, platform-polarized individuals may not revert to a more moderate state merely because exposure to the platform or content is reduced (hysteresis, see Fig. 1B). This applies not just to individuals, but likewise to the structure of the network. In the case of polarization, for example, the network of co-partisan individuals one follows could be heavily shaped by algorithmic recommendations, yet these changes are not undone by temporarily down-ranking co-partisan content in one’s own feed (Spillover, see Fig. 1D).
The experiments that change some aspects of the news feed algorithm also clearly violate SUTVA as the treatment outcome depends on the status of others. To reduce co-partisan content, one study algorithmically down-ranked such content in treated users’ feeds  (?). What a user ultimately sees is, however, also dependent on a candidate set of posts shared by their connections, which is implicitly shaped by how algorithmic curation impacts their feeds. As the study was only carried out on a small portion of the U.S. population, whatever co-partisan and cross-cutting items participants saw were determined in large part by individuals who were exposed to the unmodified algorithm. This directly parallels the example of altering the load on specific power stations, without accounting for compensatory dynamics between individual units.
The most recent Facebook deactivation study by Allcott et al.  (?), by contrast, exhibits inferential challenges from hysteresis. To the extent that users’ voting decisions were influenced by Facebook, opinions on the incumbent candidate (and perhaps the challenger, if well known) would have been formed over years prior to the election and experiment. There is no empirical reason to expect that withdrawal from Facebook has a symmetric effect to being exposed to the algorithm (likely for much longer time periods). The observed point-estimated magnitude of the vote shift being on the scale of what might tip an election (2.6%) is therefore striking, as even a zero percent change could be entirely consistent with preferences accrued over many years. The fact that the change was nonetheless non-significant (after adjusting for multiple comparisons) highlights an additional problem of acquiring sufficient statistical power in individual-level experiments to identify effects that are meaningful at the system level.
Taken together, while the Meta collaboration studies differ considerably in implementation and focus, they all similarly fail to account for the complexity inherent in online social networks. And while the authors describe some of these challenges as limitations, they are more accurately viewed as fundamental barriers to meaningful causal inference. (?)
Unactionable Research
Even if we could conclude with some confidence that collective social-media use negatively impacted some outcome we care about—for example, suppose we found that it increased affective polarization—what would we do with this knowledge? Normally, causal research guides policy-making by helping identify the causes of observed effects towards proposing interventions, or by measuring the relative effect of proposed interventions to facilitate decision-making regarding which policies or changes to adopt.
If we were to learn that social media use generally increases polarization, it could lead to the insight that decreasing or eliminating social media use might reduce affective polarization. However, social media has become digital public infrastructure with an array of effects, some good and some bad and its elimination is only thinkable for specific, vulnerable groups, such as children (a policy decision that indeed was recently made in Australia). Embracing rather than erasing heterogeneity across users is an important lever for developing interventions that can be applied to address specific harms. For instance, rather than asking whether or not social media as a whole causes decreasing well-being on average, one could instead ask if engagement-based ranking leads to unhealthy overuse.
It follows that rather than focus on the net effect, actionable research on social media should seek to identify realistic interventions concerning specific affordances of platforms that could actually be improved  (?). This could include platform interventions, such as bridging or down-ranking of toxic content, or user-side interventions such as cross-partisan dialogue  (?). Some of the Facebook studies  (?, ?) actually tested realistic interventions in the form of altered ranking algorithms, but did not assess their collective outcomes at the group-level. We encourage the research community to shift focus in this direction, in combination with the appropriate methods. In short, evaluating interventions has three key benefits over evaluating the effects of social media as a whole: it involves a coherent counterfactual and is thus a conceptually meaningful estimand; it is practically possible to study, because smaller treatments can be put in place for longer than e.g., deactivation studies; and it can guide the design of policy interventions.
In complex systems, as we described before, hysteresis and other time-dependent phenomena, such as feedback loops, play an important role. It is therefore of importance to note that the effects of interventions will depend on time and/or the state of other variables at that point in time  (?). Consequently, their effects need to be continuously evaluated, as is likely the case for commercial indicators, such as dwell-time and clicks, already (?). Interventions might be demographically or locality limited and applied for a finite time-period, whereas others may need global and persistent changes. If these interventions aim at collective outcomes, the methods for assessing their causal effects must also account for other properties of complex systems, namely that outcomes are not simply the sum of their parts and that interventions spill over through social networks.
Adding the Complex Systems Perspective
The challenges of studying social media just described are serious, and our discussion above highlights that progress forward cannot occur via solely relying on RCTs to ask unactionable questions. Fortunately, this is not uncharted territory. Conservation biologists, climate scientist, and public health researchers are rarely able to evaluate their proposed interventions using RCTs at the scale of the whole systems on which interventions are targeted. Nevertheless, they have made substantial progress in developing workable interventions for protecting ecosystems, mitigating climate change, and reducing the large-scale burden of disease. Progress in each of these areas has been heavily reliant on developing mechanistic theoretical models of the underlying processes governing the system and evaluating how interventions on that system might manifest in downstream changes. (?)
As with work in climate science and conservation, the search for interventions is often motivated by observational research. Much of the research on social media is similarly motivated by observations that suggest social media is having deleterious societal impacts. Relevant outcome variables include mental health indicators, beliefs that might undermine democracy, attitudes such as affective polarization, and behaviors such as vaccination, voting, or crime. Often these variables exhibit concerning and consistent trends, such as reductions in teen mental health, rising polarization, erosion of democratic norms, and reduced rates of vaccination. So what do we do with these patterns?
Too often, such observational findings are dismissed individually for heuristically lacking ”gold standard” status associated with RCTs. It is important to note, however, that that misunderstandings regarding the intrinsic causality of RCTs has a flip-side for observational work, where it is assumed they can solely produce associational or correlational evidence. Both misunderstandings arise from believing that a method imbues causality, when causality is instead a product of the assumptions we make about the associations we observe in our data both within and across studies  (?). Techniques of causal inference from observational data leverage naturally-occurring, rather than experimentally induced variation (?).
Of course the assumptions of any given observational study, which as with RCTs, can undermine inference. For example, many instrumental-variable studies in the field rely on the “exogenous” or “as-if random” roll out of infrastructure, this variation evaporates as technology proliferates and requires either explicitly accounting for non-random causes of roll out (e.g. socio-economic status, variation in jurisdiction) or assuming roll out is approximately random. Furthermore, instrumental-variable methods rely on the assumption that the instrument affects the outcome only through the variable of interest (the “treatment” variable), which may be difficult to justify in the complex and interconnected context of social media research (see here for a recent example of exclusion restriction violations  (?)). Finally, the interpretation of causal effects estimated through instrumental-variable analyses is limited to a specific subpopulation (the “compliers”), as the analyses focus on the estimation of local average treatment effects (LATE), which can considerably reduce the generalizability of the findings to the population of interest.
At first glance, it may seem that observational research, like RCTs, is limited by its assumptions that can undermine inference. Yet a key distinction is that a given line of observational evidence often has its own unique assumptions, such that consistency across approaches can provide some comfort that the phenomenon of interest is not merely an artifact of one particular assumption failure. In this sense, concerning observational patterns can be a guiding light towards domains that require some form of interventions. However, these large-scale and perhaps long-duration trends rarely provide firm, actionable mechanistic insight into design changes that inherently impact individuals on shorter time-scales.
Figure 2: Illustrations of the triangulation approach.
Ultimately a bridge is needed between RCTs that provides insight at the scale of individuals in discrete moments, and observational approaches that reveal trends at larger time-scales. Other complexity minded disciplines have built such bridges using an empirically-grounded theory- and model-driven paradigm that aims at understanding specific mechanisms. Such an approach of triangulating experimental, observational and theoretical work (Fig. 2), which we feel is necessary for the field to move forward, stands in contrast to relying on single-shot experiments to generate broad statements about platforms’ overall impacts. Although uncommon in the social sciences, this paradigm is not out of reach. Theory and mechanistic models do not serve to replace empirical research, but rather augment it by bridging gaps across scales of organizational complexity – from individual behavior to societal-level phenomena. If observational approaches to causal inference reveal a trend, theoretical approaches can be used to propose how such trends might arise from mechanisms at play. In turn, the mechanistic assumptions of the multi-scale model can be identified and confirmed at individual scales with formal experiments, providing confidence that the model is able to capture leading-order features governing dynamics across scales. Computational methods, such as agent-based simulations or interactive online experiments furthermore allow for bridging the methodological gaps.
The spread of misleading information online provides a salient example of how triangulation across observational and experimental research can be facilitated by complexity-minded theory. Given observations that misinformation proliferates online, a variety of interventions have been proposed, studied, experimentally validated in RCTs  (?). However, despite the success in RCTs and even convincing some platforms to adopt these approaches, estimating their population-level effect remains is not straightforward as the dynamics of misinformation spreading present a typical example of a complex, non-linear system  (?).
Yet misinformation spreads in a manner that can be captured by models of contagious processes  (?). Beyond mere theory, data can be used to constrain parameters of contagion models, providing a tight fit to observed dynamics and allowing for evaluation of counterfactuals in which various policies are implemented  (?, ?). Contagion models can be explicitly made consistent with evidence from RCTs, by hard-coding the observed reductions in sharing behavior from experiments. Using such a model fit to data from the 2020 US elections, it was predicted that despite the nominally large effect size of the “community notes” program which relies on appending crowd-sourced fact-checks  (?), changes in exposure would be minimal—as large amounts of spread by influential accounts can occur prior to labeling and for unlabeled posts. Observational investigations provided confirmation of these model implications, confirming that even as community notes were indeed too slow to significantly reduce engagement with misinformation  (?).
These mechanistic models can often be used to evaluate multiple potential interventions in a common framework. Here, for example, the model indicates that dynamics of information spread on platforms will be highly determined by the actions of influential accounts  (?). While content creators have disproportionate agency in shaping the content that is created and propagated, which is sometimes cited as an argument against the relative importance of social media platforms, a complex systems view recognizes their role in dependence of specific incentive structures that are set by the platforms  (?). For example, by using economic models, as content creation on platforms is labor  (?). And this suggestion that influence incentives provide a uniquely large lever for intervention is indicated on longer time-horizons by observational research by Frimer et al., which measured a 23% increase in incivility among US politicians and attributed it partly to learning from positive feedback for uncivil tweets  (?). As interventions at the creator level have been comparatively understudied relative to user-specific interventions, this suggests a fruitful avenue for future research. For example, we might ask what can break feedback loops between creators and their audiences that may be driving influencers towards extremism.
The critical thing is that we can either, as a field, opt to continue scratching our heads over conflicting evidence or attempt to bring it all together into a common framework. Recognizing large-scale patterns that are consistent across observational approaches which vary in their assumptions can guide us towards where interventions may be needed. RCTs can help us understand proximate, small-scale effects that in turn can fuel theory linking observations and intervention. In triangulating across these modalities of inference, research on social media can follow in the footsteps of other disciplines that have embraced complexity to develop workable interventions in complex systems (Fig. 2.
Discussion
A considerable body of research on the causal impacts of social media has chased an unanswerable, and unactionable question of whether the world would be, or would have been, better off without social media. Disproportionate weight has been given to RCTs which, in failing to account for complexity, cannot yield the desired insight into impacts of platforms at societal scales. Trying to reconcile evidence from RCTs with either proximal patterns of content exposure or more distal patterns that emerge at societal scales, such as polarization, is a recipe for failing to reach consensus and inaction.
More generally, it is worth reflecting on our inferential goals when establishing links between platforms and harms. Such quantities are likely to be time-varying, hysteretic, and heterogeneous such that it is unlikely that we could come up with a single estimate of effects on any given domain of concern—much less across domains. Even if we could, such an estimate could do little more than motivate an attempt to extricate social media from society—an intervention that is difficult to execute and could very well do more harm than good. Instead, the key value in any such apparent causal links, however tenuous, is a recognition that social media can impact large-scale phenomena of interest. This, in turn, motivates the search for design principles and regulatory action that can reduce risks and enhance benefits.
Perhaps the first question we can ask is whether we need a causal understanding of larger-scale effects to warrant intervention. Normative and legal justifications for altering the distribution of content may in many cases be sufficient. Researchers may be able to best support such interventions by simply documenting failures of platforms preventing the distribution of illegal content of their own accord. The Digital Services Act of the European Commission is a step towards enabling researchers to do so. Platform interventions could go beyond content moderation and build on our existing understanding of how algorithmic design choice shapes content amplification. At present these choices are decided based on business priorities, and there is little reason they should not, or cannot reflect widely held prosocial goals. Moreover, as the content shared on a platform is disproportionately shaped by the incentives that govern content creation, understanding the behavior and dynamics of content creators is likely to be a particularly productive line of research.
When more firm causal links across scales of organizational complexity are required, it is important to recognize that no approach in social media research represents the “gold standard” possible in other scientific domains. As a result, we should not anticipate that any given modality of inference (e.g., RCTs, observation, theory) will be sufficient to identify phenomena and propose interventions. Instead, progress will require embracing complexity and relying on a triangulation across lines of evidence with varying assumption, leveraging computational and mathematical models to link and infer dynamics across time and scale.
We believe this is not only a possible path forward for the field, but a necessary one. The alternative is to endlessly fail to reconcile conflicting evidence, chasing the ghosts of overlooked assumptions and unanswerable questions. Doing so not only fails to yield actionable insight, but undermines any motivations by stakeholders to search for interventions in the first place. Given this, the community should set aside notions of rote causal impacts and move towards more actionable, complexity minded, and informative social media research.
Declarations
  • Funding. SL and PLS acknowledges financial support from the Volkswagen Foundation (grant “Reclaiming individual autonomy and democratic discourse online: How to rebalance human and algorithmic decision making”) and the European Commission (Horizon 2020 grant 101094752 SoMe4Dem). SL also acknowledges funding from the European Research Council (ERC Advanced Grant 101020961 PRODEMINFO), the Humboldt Foundation through a research award. SL also receives funding from UK Research and Innovation (through EU Horizon replacement funding grant number 10049415). AO is funded by the Medical Research Council (MC/UU/0030/13 and MR/X028801/1), a UK Research and Innovation Future Leader’s Fellowship (MR/X034925/1) and the Jacobs Foundation. J-B.C: A grant from Templeton World Charities (TWCF-2023-32581) supported this research. L.O.: European Commission (Horizon 2020 grant 101094752 SoMe4Dem)
  • Competing interests. SL also receives funding from Jigsaw (a technology incubator created by Google). SL, PLS and AO also work with the Joint Research Centre of the European Union, both in a paid and unpaid capacity. J.B.-C. has engaged in paid consulting for the United Nations.